# Stochastic Kinetics of mRNA Molecules in a General Transcription Model

### v3

@@ 1 Introduction
Furthermore, when working with CMEs, most previous studies consider only the steady-state solution by manually setting all time derivatives to zero. In such cases, the existence of equilibrium is presumed, which is not necessarily valid. Also, little is known about temporal evolution of chemical reaction kinetics, since convergence to equilibrium and the rate of approaching equilibrium both remain obscure. Also, the dynamical behavior of the chemical reaction kinetics remains poorly characterized. Both the conditions guaranteeing convergence to equilibrium and the corresponding relaxation times are not well understood.
@@ 2.2 Chemical Master Equation and its Solution
In (4) and (5), $\Omega_{k}:=\{(t_{1},t_{2},\cdots,t_{k})\mid 0\leq t_{1}\leq t_{2}\leq\cdots\leq t_{k}\leq t\}\subseteq\mathbb{R}^{k}$; #\# denotes the cardinality of a set; $\alpha_{s}:=\exp(-\delta t+\delta t_{s})$; $I^{c}$ denotes the complement of $I$ with respect to $\{1,\cdots,k\}$; $\{\bm{e}_{s}\}_{s=1}^{N}$ are standard basis vectors in $\mathbb{R}^{N}$ where $\bm{e}_{s}$ is the column vector with a $1$ in the $s$-th position and 0 elsewhere. The kernel function is defined as follows: $K(t;t_{1},\cdots,t_{k}):=\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\cdots D_{1}\mathrm{e}^{D_{0}(t_{k}-t_{k-1})}D_{1}\mathrm{e}^{D_{0}(t-t_{k})}$ for $k\geq 3$, $K(t;t_{1}):=\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t-t_{1})}$, and $K(t;t_{1},t_{2}):=\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\mathrm{e}^{D_{0}(t-t_{2})}$. The exponential of a square matrix, say $C$, is defined as $\mathrm{e}^{C}:=\sum_{k=0}^{\infty}C^{k}/k!$ (see Chapter $9$ of [10]). We note that the summation over $I$ in (4) ranges over all unordered $m$-element subsets without repetition, and, when $I^{c}$ is the empty set, we take the corresponding term in the product as $1$. Detailed derivation is in Appendix, Section A.
@@ 2.3 Binomial Moment Method
$B_{m}$ $=\frac{1}{m\delta}\bm{\pi}^{\top}D_{1}\prod_{k=1}^{m-1}\left[\left(k\delta\bm{I}_{N}-D\right)^{-1}D_{1}\right]\bm{1},\quad m\geq 2,\;\;\text{and}\;\;B_{1}=\frac{1}{\delta}\bm{\pi}^{\top}D_{1}\bm{1},$ (8)
@@ 3.2 Special Case I: The Telegraph Model
$B_{m}$ $=\frac{1}{m!}\left(\frac{\lambda}{\delta}\right)^{m}\frac{(k_{2}/\delta)^{(m)}}{(k_{1}/\delta+k_{2}/\delta)^{(m)}},$ (11)
@@
$\mathcal{P}_{n}$ $=\frac{1}{n!}\frac{\lambda^{n}(k_{2}/\delta)^{(n)}}{\delta^{n}(k_{1}/\delta+k_{2}/\delta)^{(n)}}\;\mathllap{{}_{1}}F_{1}\left(\frac{k_{2}}{\delta}+n,\frac{k_{1}+k_{2}}{\delta}+n;-\frac{\lambda}{\delta}\right),$ (12)
@@ 3.3 Special Case II: Markovian Models Under Renewal Condition
Recall that a renewal process is a counting process whose interarrival times are independent and identically distributed according to a given distribution with probability mass function $f(t)$. Note that $f(t)$ is assumed to be integrable in most cases. Previous studies predominantly focus on the case where production of mRNA molecules becomes a renewal process, and queueing theory is introduced to obtain powerful results in a relatively general framework [20, 26, 40, 43]. In (1), production of mRNA molecules becomes a renewal process if and only if the following renewal condition [43] holds: When $\sum_{j=1}^{N}b_{i,j}\neq 0$, $\frac{b_{i,j}}{\sum_{j=1}^{N}b_{i,j}}$ is independent of $i$. Assuming $D_{1}\neq\bm{0}_{N\times N}$, this renewal condition is equivalent to $\text{rank}\,D_{1}=1$.
We specifically term (1) as a Markovian model under renewal condition when the aforementioned renewal condition is satisfied. Note that the Telegraph model discussed in the previous subsection can also be incorporated into this framework; however, we treat it separately because its results permit further simplification. In general, the Markovian model under renewal condition can be interpreted alternatively as a queueing system $GI/M/\infty$, with probability mass function of interarrival times being [43]
$f(t)=\bm{\kappa}^{\top}\mathrm{e}^{D_{0}t}D_{1}\bm{1},$ (13)
@@ 3.4 Numerical Analysis
On the other hand, floating point error is introduced in both (8) and (7). Computing binomial moments via (8) is relatively accurate and stable, because inversion of diagonally dominant matrix, matrix-matrix multiplication and matrix-vector multiplication are all controllable operations [10]. However, floating point error can become extremely severe when reconstructing probability mass function based on (7). Since binomial moments may rapidly drop below the threshold of machine precision while the combinatorial coefficients can grow very large, multiplying a small binomial moment by a large combinatorial coefficient suffers from precision loss. Thereby, particular care must be taken in practical implementations. The essentially same problem has been observed and carefully analyzed in [13]. In the Jupyter Notebooks provided in the Supplementary Material, we compute terms $\binom{m}{n}B_{m}$ in (7) by first performing logarithmic-scale operations (use built-in function math.lgamma(n+1) to compute $\ln{(n!)}$ and then exponentiating the output. Our experience shows that this approach is much more stable than direct computation using math.comb.
@@ B.2 Take Temporal Limit
$=$ $\frac{1}{m\delta}\bm{\pi}^{\top}D_{1}\prod_{k=1}^{m-1}\left[\left(k\delta\bm{I}_{N}-D\right)^{-1}D_{1}\right]\bm{1},\quad m\geq 2;$
@@ Appendix C Upper Bound for Binomial Moments (9)
$B_{m}$ $=\frac{1}{m\delta}\bm{\pi}^{T}D_{1}\left(\delta\bm{I}_{N}-D\right)^{-1}D_{1}\left(2\delta\bm{I}_{N}-D\right)^{-1}\cdots D_{1}\left[(m-1)\delta\bm{I}_{N}-D\right]^{-1}D_{1}\bm{1}$ (S.22)

### v4

@@ 1 Introduction
Furthermore, when working with CMEs, most previous studies consider only the steady-state solution by manually setting all time derivatives to zero. In such cases, the existence of equilibrium is presumed, which is not necessarily valid. Also, the dynamical behavior of the chemical reaction kinetics remains poorly characterized. Both the conditions guaranteeing convergence to equilibrium and the corresponding relaxation times are not well understood.
@@ 2.2 Chemical Master Equation and its Solution
In (4) and (5), $\Omega_{k}:=\{(t_{1},t_{2},\cdots,t_{k})\mid 0\leq t_{1}\leq t_{2}\leq\cdots\leq t_{k}\leq t\}\subseteq\mathbb{R}^{k}$; #\# denotes the cardinality of a set; $\alpha_{s}:=\exp(-\delta t+\delta t_{s})$; $I^{c}$ denotes the complement of $I$ with respect to $\{1,\cdots,k\}$; $\{\bm{e}_{s}\}_{s=1}^{N}$ are standard basis vectors in $\mathbb{R}^{N}$ where $\bm{e}_{s}$ is the column vector with a $1$ in the $s$-th position and 0 elsewhere. The kernel function is defined as follows: $K(t;t_{1}):=\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t-t_{1})}$, $K(t;t_{1},t_{2}):=\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\mathrm{e}^{D_{0}(t-t_{2})}$, and
$K(t;t_{1},\cdots,t_{k}):=\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\cdots D_{1}\mathrm{e}^{D_{0}(t_{k}-t_{k-1})}D_{1}\mathrm{e}^{D_{0}(t-t_{k})}$ for $k\geq 3$. The exponential of a square matrix, say $C$, is defined as $\mathrm{e}^{C}:=\sum_{k=0}^{\infty}C^{k}/k!$ (see Chapter $9$ of [10]). We note that the summation over $I$ in (4) ranges over all unordered $m$-element subsets without repetition, and, when $I^{c}$ is the empty set, we take the corresponding term in the product as $1$. Detailed derivation is in Appendix, Section A.
@@ 2.3 Binomial Moment Method
$B_{1}$ $=\frac{1}{\delta}\bm{\pi}^{\top}D_{1}\bm{1},$ (8)
$B_{m}$ $=\frac{1}{m\delta}\bm{\pi}^{\top}D_{1}\left(\delta\bm{I}_{N}-D\right)^{-1}D_{1}\left(2\delta\bm{I}_{N}-D\right)^{-1}\cdots D_{1}\left[(m-1)\delta\bm{I}_{N}-D\right]^{-1}D_{1}\bm{1},\;m\geq 2,$
@@ 3.2 Special Case I: The Telegraph Model
$B_{m}$ $=\frac{1}{m!}\left(\frac{\lambda}{\delta}\right)^{m}\frac{(k_{2}/\delta)^{(m)}}{(k_{1}/\delta+k_{2}/\delta)^{(m)}},\;m\in\mathbb{N},$ (11)
@@
$\mathcal{P}_{n}$ $=\frac{1}{n!}\frac{\lambda^{n}(k_{2}/\delta)^{(n)}}{\delta^{n}(k_{1}/\delta+k_{2}/\delta)^{(n)}}\;\mathllap{{}_{1}}F_{1}\left(\frac{k_{2}}{\delta}+n,\frac{k_{1}+k_{2}}{\delta}+n;-\frac{\lambda}{\delta}\right),\;n\in\mathbb{N},$ (12)
@@ 3.3 Special Case II: Markovian Models Under Renewal Condition
Recall that a renewal process is a counting process whose interarrival times are independent and identically distributed according to a given distribution with probability density function $f(t)$. Note that $f(t)$ is assumed to be integrable in most cases. Previous studies predominantly focus on the case where production of mRNA molecules becomes a renewal process, and queueing theory is introduced to obtain powerful results in a relatively general framework [20, 26, 40, 43]. In (1), production of mRNA molecules becomes a renewal process if and only if the following renewal condition [43] holds: When $\sum_{j=1}^{N}b_{i,j}\neq 0$, $\frac{b_{i,j}}{\sum_{j=1}^{N}b_{i,j}}$ is independent of $i$. Assuming $D_{1}\neq\bm{0}_{N\times N}$, this renewal condition is equivalent to $\text{rank}\,D_{1}=1$.
We specifically term (1) as a Markovian model under renewal condition when the aforementioned renewal condition is satisfied. Note that the Telegraph model discussed in the previous subsection can also be incorporated into this framework; however, we treat it separately because its results permit further simplification. In general, the Markovian model under renewal condition can be interpreted alternatively as a queueing system $GI/M/\infty$, with probability density function of interarrival times being [43]
$f(t)=\bm{\kappa}^{\top}\mathrm{e}^{D_{0}t}D_{1}\bm{1},\;t>0,$ (13)
@@ 3.4 Numerical Analysis
On the other hand, floating point error is introduced in both (8) and (7). Computing binomial moments via (8) is relatively accurate and stable, because inversion of diagonally dominant matrix, matrix-matrix multiplication and matrix-vector multiplication are all controllable operations [10]. However, floating point error can become extremely severe when reconstructing probability mass function based on (7). Since binomial moments may rapidly drop below the threshold of machine precision while the combinatorial coefficients can grow very large, multiplying a small binomial moment by a large combinatorial coefficient suffers from precision loss. Thereby, particular care must be taken in practical implementations. The essentially same problem has been observed and carefully analyzed in [13]. In the Jupyter Notebooks provided in the Supplementary Material, we compute terms $\binom{m}{n}B_{m}$ in (7) by first performing logarithmic-scale operations (use built-in function math.lgamma(n+1) to compute $\ln{(n!)}$ ) and then exponentiating the output. Our experience shows that this approach is much more stable than direct computation using math.comb.
@@ B.2 Take Temporal Limit
$=$ $\frac{1}{m\delta}\bm{\pi}^{\top}D_{1}\left(\delta\bm{I}_{N}-D\right)^{-1}D_{1}\left(2\delta\bm{I}_{N}-D\right)^{-1}\cdots D_{1}\left[(m-1)\delta\bm{I}_{N}-D\right]^{-1}D_{1}\bm{1},\;m\geq 2;$
@@ Appendix C Upper Bound for Binomial Moments (9)
$B_{m}$ $=\frac{1}{m\delta}\bm{\pi}^{\top}D_{1}\left(\delta\bm{I}_{N}-D\right)^{-1}D_{1}\left(2\delta\bm{I}_{N}-D\right)^{-1}\cdots D_{1}\left[(m-1)\delta\bm{I}_{N}-D\right]^{-1}D_{1}\bm{1}$ (S.22)